OpenStudy (loser66):

A-72kg trampoline artist jumps vertically upward from the top of a platform with a speed of 4.5m/s. a) How fast is he going as he lands on the trampoline 2.0m below? b) If the trampoline behaves like a spring of spring constant 5.8x10^4N/m, how far does he depress it? Please help

OpenStudy (loser66):

for part a) I use \(KE_{initial}+ PE_{initial}= KE_{final}+KE_{final}\) \[\frac{1}{2}mv_i^2+mgh_0 = \frac{1}{2}mv_f^2+mgh_f\] \(v_f= 7.7m/s\) but I don't get part b which answer is -0.28 m

OpenStudy (loser66):

@ybarrap hey, help me, friend

OpenStudy (ybarrap):

It's asking you to use Hooke's Law. Do you know it?

OpenStudy (loser66):

yes, F= -kx

OpenStudy (loser66):

the net is soooo bad here

OpenStudy (ybarrap):

(I know, I think the server is slow today) I think this is the approach: Hooke's Law is F=kx, where k is a constant and x is the displacement away from rest: Using Newton's 2nd Law: $$ F(x)=ma(x)=-mg-(-kx)=-mg+kx $$ At maximum displacement, \(x_{max}\), velocity is zero and not upward or downward acceleration, i.e. a=0. So, $$ F(x_{max})=ma(x_{max})=0=-mg+kx_{max} $$ Now, let k=5.8x10^4N/m and solve for \(x_{max}\), the maximum displacement.

OpenStudy (ybarrap):

It says I'm typing a reply but I'm not lol

OpenStudy (anonymous):

Anyone tried using a graphics tablet to work with maths software ?